04:20
Did another temperature cycle (118 K -> 125 K -> 95 K -> 116 K) trying to improve the pressures at the target and the detector count rate.
Rates are still mostly below 100 counts (Ex rate) and the pressures are still fluctuating by about an order of magnitude (E4: between 1e-9 and 1e-8 mbar). |